# Skyfall-GS: Synthesizing Immersive 3D Urban Scenes from Satellite Imagery

### v1

@@ 4.1 Comparisons with Baselines
Table 3: Ablation on the reconstruction stage. Appearance modeling secure convergence. Opacity regularization and depth supervision enhance visual fidelity and geometric accuracy.
@@
Supervision
$\text{FID}_{\text{CLIP}}\,\downarrow$ CMMD $\downarrow$
Failed Failed
41.90 2.45
39.95 2.40
38.01 2.31
Table 4: Ablation on the synthesis stage. Multiple samples and curriculum learning both significantly improve rendering quality, achieving the best scores in all metrics.
Multiple
Samples
Curriculum
Learning
$\text{FID}_{\text{CLIP}}\,\downarrow$ CMMD $\downarrow$
34.11 3.19
33.79 3.36
28.35 2.88
Figure 8: Satellite-view training and IDU refinement ablation. The opacity regularization reduces floating artifacts and yields denser reconstructions. The pseudo-camera depth supervision improves geometry in texture-less areas like rooftops and roads. The multiple diffusion samples per view enhance texture consistency and geometry. The curriculum learning progressively introduces challenging views, significantly improving geometric coherence in previously occluded regions.
4.2 Ablation Studies
We conduct ablation studies on the JAX_068 AOI data.
Ablation on the reconstruction stage. We ablate appearance modeling, opacity regularization, and pseudo-camera depth supervision (see Table˜4 and Figure 8). For this ablation, we evaluate at higher elevation angles to assess the quality of renders during the IDU process, rather than testing the final low-angle performance. Appearance modeling is crucial for multi-date convergence, opacity regularization removes floating artifacts (Figure˜8(a)), and depth supervision flattens planar regions (Figure˜8(b)). Together, they yield the lowest $\text{FID}_{\text{CLIP}}$/CMMD scores.
Ablation on the synthesis stage. We isolate two factors: multi-sample diffusion and curriculum view progression. As Figure˜8(c-d) shows, multi-sampling smooths textures, while the curriculum (vs. random views) restores geometry in occluded areas. Table˜4 confirms these gains.
@@ 5 Conclusion
Skyfall-GS synthesizes real-time, immersive 3D urban scenes from multi-view satellite imagery, using 3D Gaussian Splatting and text-to-image diffusion models in a curriculum-based iterative refinement approach. Our method surpasses existing methods like Sat-NeRF, , EOGS, CityDreamer, and GaussianCity, effectively addressing challenges such as limited parallax, illumination variations, and occlusions. Future work includes scaling to larger environments and dynamic scenes.
Limitations. Our method requires significant computational resources, primarily due to the refinement process. The approach produces over-smoothed textures at extreme street-level perspectives. Future work will focus on reducing computational overhead, improving street-level detail, and extending the approach with robust geometric validation.
Ethics Statement
This work included a small-scale user study where anonymous participants were asked to compare our results with baselines through an online survey. No personally identifiable information was collected, and all responses were stored anonymously. Participation was entirely voluntary, and no risks were posed to participants. The study did not require institutional review board (IRB) approval under our institution’s policies, as it involved only anonymous survey responses with minimal risk.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. Implementation details of our reconstruction and synthesis pipeline are provided in Section 3, including the architecture, loss functions, and optimization objectives. All hyperparameters, training schedules, and regularization terms are described in Section˜3 and Section˜A.1. Details of datasets, splits, and evaluation protocols are described in Section˜4 and Section˜A.2, with clear references to the publicly available DFC2019 dataset and GoogleEarth dataset. Details of the user study are described in Section˜A.2. We will release our source code upon acceptance to further support transparency and reproducibility.
@@ Acknowledgements
Appendix A Appendix
This supplementary material provides additional details that complement our main paper. We include:
@@ Appendix A Appendix
    Implementation Details: Pseudo camera depth supervision strategy, 3DGS reconstruction parameters for satellite imagery, and FlowEdit-based refinement process.
@@
    Experimental Information: Dataset details with training image counts and geographical coordinates for each Area of Interest (AOI), user study methodology, and evaluation protocol.
@@
    Additional Qualitative Results: Extended visual comparisons with state-of-the-art methods and results on four additional AOIs of the DFC2019 dataset demonstrating the robustness of our approach.
Additionally, we provide an interactive HTML visualization (available in the folder, main.html) that allows readers to explore our video results and compare reconstructions across different viewing conditions and scenes. This visualization enables direct comparison of our method’s geometric accuracy, spatial alignment, and overall perceptual quality against baseline approaches and Google Earth Studio reference video.
We also provide example datasets via Zenodo, which can be accessed at this URL. However, due to storage limitations, we only provide training data for an AOI as an example. We plan to release the complete dataset upon acceptance.
A.1 Implementation Details
Codebase. Our method extends the Mip-Splatting (Yu et al., 2024) codebase with custom modules for satellite imagery processing and our curriculum-based IDU refinement pipeline.
Pseudo camera depth supervision. We sample cameras with varied azimuths and decreasing elevations, using random per-image embeddings. MoGe (Wang et al., 2024a) provides scale-invariant depth estimation. We sample 24 views every 10 iterations, with look-at points $(x,y,z)$, where $x,y\sim\mathcal{N}(0,128)$ and $z=0$, as illustrated in Figure˜9. Camera azimuths are uniformly sampled between 0 and $2\pi$, while elevation angles and radii linearly decrease from $80\degree$ to $45\degree$ and 300 to 250 units, respectively. Rendered RGB images ($I_{\text{RGB}}$) are $1024\times 1024$ pixels. We illustrate the 3DGS rendered RGB image $I_{\text{RGB}}$, scale-invariant depth $D_{\text{est}}$ estimated by MoGe (Wang et al., 2024a) and depth from 3DGS $D_{\text{GS}}$ in Figure˜10.
Figure 9: The sampling strategy of pseudo camera. In this example, we sample 240 points using the strategy.
@@ A.1 Implementation Details
Figure 10: Pseudo-cam Depth Supervision. We use MoGe (Li et al., 2024c) to estimate the scale-invariant depth $D_{\text{est}}$ from the rendered RGB image $I_{\text{RGB}}$. The rightmost figures show the rasterized depth $D_{\text{GS}}$ from 3DGS.
@@
3DGS reconstruction from satellite imagery. Our satellite-view optimization process runs for 30,000 iterations, with densification enabled between iterations 1,000 and 21,000. We modify several key parameters in the standard 3DGS implementation to address satellite imagery’s unique challenges. First, to prevent undesirable Gaussian elongation artifacts common with overhead views, we reduce the scaling learning rate from 0.005 to 0.001. Second, we address sparsity issues of Gaussian points in close-up renderings by lowering the densification gradient threshold from 0.002 to 0.001, ensuring sufficient detail when viewed from ground level. Furthermore, we implement pruning of Gaussians with maximum covariance exceeding 20 to eliminate floating artifacts. The loss function weights are set to $\lambda_{\text{D-SSIM}}=0.2$, $\lambda_{\text{op}}=10$, and $\lambda_{\text{depth}}=0.5$ for optimal reconstruction quality. For appearance modeling, we adopt the architecture from WildGaussians (Kulhanek et al., 2024), implementing an appearance MLP with 2 hidden layers (128 neurons each) and ReLU activation functions. The per-image and per-Gaussian embedding dimensions are set to 32 and 24 respectively, with learning rates of 0.001, 0.005, and 0.0005 for per-image embeddings $e_{j}$, per-Gaussian embeddings $g_{i}$, and the appearance MLP $f$, respectively. The complete satellite-view training requires approximately 1 hour on a single NVIDIA RTX A6000 GPU.
@@
FlowEdit-based refinement. We set FlowEdit noise parameters $n_{\text{min}}=4$ and $n_{\text{max}}=10$ to balance artifact removal with detail preservation. Our source prompt (“Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts.”) characterizes initial renders, while the target prompt (“Clear satellite image of an urban area with sharp buildings, smooth edges, natural lighting, and well-defined textures.”) guides refinement. These parameters were determined through experimentation, with lower noise values preserving more original structure but removing fewer artifacts, and higher values creating more significant changes but potentially altering underlying geometry. All other FlowEdit parameters use default values.
@@
Curriculum-based refinement details. Our IDU process comprises $N_{e}=5$ episodes of 10,000 iterations each, with densification through iteration 9,000. At the start of IDU, we randomly select and fix a single per-image appearance embedding $e_{j}$. Opacity regularization is disabled during IDU, as our curriculum naturally mitigates floating artifacts through multi-view consistency, enabling Gaussians to retain variable opacities beneficial for semi-transparent structures (Kerbl et al., 2023). For DFC2019 (Le Saux et al., 2019) dataset, we utilize $N_{p}=9$ look-at points in a $3\times 3$ grid (512 units wide, centered at origin), with $N_{v}=6$ cameras per point and $N_{s}=2$ samples per view. Camera elevations decrease from $85^{\circ}$ to $45^{\circ}$ and radii from 300 to 250 units across episodes. For GoogleEarth (Xie et al., 2024) dataset, we utilize $N_{p}=16$ look-at point at origin, with $N_{v}=6$ cameras per point and $N_{s}=2$ samples per view. Camera elevations decrease from $85^{\circ}$ to $45^{\circ}$ and radius is fixed 600-unit across episodes. All training images are rendered at $2048\times 2048$ resolution. Our training strategy samples 75% from refined images and 25% from original satellite images, this sampling strategy makes sure that the final 3DGS scene faithfully follows the semantic and layout in the input satellite imagery. The complete synthesizing stage requires approximately 6 hours on a single NVIDIA RTX A6000 GPU.
A.2 Experiments
@@
DFC2019 (Le Saux et al., 2019) dataset details. The number of training images and geographical coordinates for each AOI is provided in Table˜5. We also include four additional AOIs from Jacksonville to demonstrate our method’s robustness across varying scene characteristics. The number of training images and geographical coordinates for these additional AOIs is provided in Table˜6. These additional AOIs feature distinct characteristics: one contains a city hall building (JAX_164), another includes an American football stadium (JAX_175), while the remaining two exhibit other notable urban features (JAX_168 and JAX_264).
@@
GoogleEarth (Xie et al., 2024) dataset details. The GoogleEarth dataset, introduced by CityDreamer (Xie et al., 2024), contains semantic maps, height fields and renders from Google Earth Studio (Google, 2024) of New York City. This dataset is used to train the generative model in CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b). We pick four AOIs which contain diverse city elements, including complex architectures (004), squares (010), resident area (219) and riverside (336). However, original GES renders provided in GoogleEarth dataset are rendered from a lower elevation angle, which is not similar to satellite imagery. Therefore, for each AOI, we render 60 images from GES using an orbit trajectory with $80^{\circ}$ of elevation angle and 2219 of radius. These new renders serve as the input of our methods. The AOI ID, geographical coordinates, and the number of input images are detailed in Table˜7.
@@
Table 5: Number of training images and geographical coordinate per Area of Interest (AOI). These AOIs correspond to standard evaluation scenarios established by previous works, ensuring consistent and fair comparisons with existing baselines (e.g., Sat-NeRF (Marí et al., 2022)).
@@
Table 7: Number of training images and geographical coordinate per Area of Interest (AOI). We pick 4 AOIs from the GoogleEarth (Xie et al., 2024) dataset, ensuring fair comparisons with existing baselines (e.g., CityDreamer  (Xie et al., 2024) and GaussianCity (Xie et al., 2025b))
@@
Figure 11: Additional qualitative comparison on the DFC2019 dataset with Sat-NeRF (Marí et al., 2022), Mip-Splatting (Yu et al., 2024), CoR-GS (Zhang et al., 2024b), and EOGS (Savant Aira et al., 2025). Our method significantly outperforms baseline approaches in both geometric accuracy and texture quality when rendering low-altitude novel views. Note the superior building geometry, facade details, and reduced floating artifacts in our final result.
Figure 12: Additional qualitative comparison on the GoogleEarth dataset with CityDreamer (Xie et al., 2024), GaussianCity (Xie et al., 2025b), and CoR-GS (Zhang et al., 2024b). Our method is able to synthesize texture and geometry that is closer to the reference GES render.
User study details. We asked participants three specific questions and instructed them to select one video that best addressed each question:

### v2

@@ 4.2 Ablation Studies
Table 3: Ablation on the reconstruction stage. Appearance modeling secures convergence. Opacity regularization and depth supervision enhance visual fidelity and geometric accuracy.
Components Perceptual Metrics Geometric Metrics
@@
Sup.
[…]
Failed Failed Failed Failed
41.90 2.450 3.542 5.218
39.95 2.395 2.980 4.527
38.01 2.307 2.250 3.483
Table 4: Ablation on the synthesis stage. We evaluate sample counts ($N_{s}$), core components, and compare against baselines.
Method Variation $\text{FID}_{\text{CLIP}}\downarrow$ CMMD $\downarrow$ Time (h)
Multiple Samples ($N_{s}$)
$N_{s}=1$ 34.11 3.189 3.44
Ours ($N_{s}=2$) 28.35 2.875 6.37
$N_{s}=3$ 28.64 2.769 7.19
$N_{s}=5$ 29.17 2.677 9.80
Component Ablation
w/o Curriculum 33.79 3.361 -
w/ Context-free Pmt. 30.78 2.981 -
Replaced w/ SDEdit 64.74 4.138 -
4.2.2 Ablation on the synthesis stage.
We isolate multi-sample diffusion and curriculum view progression. $N_{s}=2$ achieves optimal visual results (Figure˜9(c)), while $N_{s}=5$ yields the lowest CMMD but requires $1.5\times$ longer training with marginal gains, so we adopt $N_{s}=2$. Curriculum scheduling outperforms random sampling in restoring occluded geometry (Figure˜9(d), Table˜4). Replacing our refinement with SDEdit [60] under the same schedule causes significant quality degradation, as noising-denoising cannot hallucinate plausible details while preserving satellite-defined geometry (Figure˜9(e), Table˜4). A context-free prompt yields negligible difference, confirming robustness to prompts, with full prompts provided in Supplementary.
4.3 Performance and Scalability
4.3.1 Training efficiency.
We evaluate runtime on the JAX_214 AOI using a single NVIDIA RTX A6000 (48GB). The full pipeline completes in $\sim$6h 45min, split across 1h 35min for reconstruction and 5h 10min for synthesis. Within each IDU episode of $\sim$1h, compute roughly split between render refinement and 3DGS update at $\sim$30min and $\sim$32min respectively, with initial rendering negligible at $\sim$4s. We consider this a reasonable trade-off, as our method entirely bypasses time- and labor-intensive physical data collection.
Figure 10: Multi-block scalability on combined JAX_214 and JAX_260 AOIs. Skyfall-GS jointly optimizes a seamless 3D scene spanning $\sim$1km $\times$ 512m from two adjacent satellite tiles (left). The zoomed inset highlights a shared building, showing no stitching artifacts and confirming scalability to multi-block environments.
4.3.2 Multi-block scalability via combined imagery.
To evaluate multi-block scalability, we combined two adjacent AOIs (JAX_214, JAX_260) that overlap along a highway into a single $\sim$1km $\times$ 512m dataset, scaling the IDU look-at grid to $6\times 3$ with 5 episodes. Training took $\sim$9 hours on a single RTX A6000, yielding $\sim$3.5M Gaussians and a $\sim$46GB memory footprint. As shown in Figure˜21, highway and buildings shared between both AOIs show no stitching artifacts and remain geometrically consistent across the boundary, confirming robust scaling to multi-block environments.
@@ 5 Conclusion
We present Skyfall-GS, the first method to synthesize real-time, immersive, and freely navigable 3D urban scenes solely from multi-view satellite imagery, requiring no human intervention or domain-specific 3D training data. By combining 3D Gaussian Splatting with open-domain diffusion priors in a curriculum-based iterative refinement strategy, our method effectively addresses long-standing challenges, including limited parallax, illumination variations, and large-scale occlusions. Extensive experiments demonstrate consistent outperformance over reconstruction and generation baselines, including Sat-NeRF, EOGS, CoR-GS, Mip-Splatting, CityDreamer, and GaussianCity. We hope this work paves the way for scalable, automated 3D urban scene creation, with promising future directions including city-wide scaling and dynamic scene modeling.
5.0.1 Limitations.
The fixed heuristic camera trajectory works well for most scenarios, but may leave blind spots in complex urban geometries, occasionally causing minor artifacts in heavily occluded regions at extreme street-level perspectives. Additionally, our framework requires off-nadir satellite views to synthesize building facades, leaving synthesis from purely nadir imagery as an open challenge.
@@ Acknowledgements
Appendix 0.A Supplementary Material
This supplementary material provides additional details complementing the main paper, organized into the following sections:
@@ Appendix 0.A Supplementary Material
    Implementation Details (Sec. 0.B): Expanded descriptions of method components including pseudo-camera depth supervision, 3DGS reconstruction parameters, and the FlowEdit-based refinement process. This section also covers system profiling (rendering efficiency, memory consumption) and mathematical validation of the RPC-to-perspective approximation.
@@
    Main Paper Experiments Detail & Results (Sec. 0.C): Comprehensive dataset statistics, the user study methodology, and the evaluation protocol. Additionally, this section provides the complete per-scene quantitative benchmarks and extended qualitative comparisons across both the DFC2019 and GoogleEarth datasets.
@@
    Additional Experiments (Sec. 0.D): Further evaluations demonstrating the robustness of our approach, including: (i) synthesis on complex irregular geometries and bridges; (ii) visualizing transient object handling; (iii) refinement prompt sensitivity analysis; (iv) multi-block scalability via combined imagery; (v) episode-vs-coverage analysis quantifying curriculum effectiveness; and (vi) structural consistency under different random seeds.
We provide an interactive HTML visualization (main.html) for exploring video results across scenes and viewing conditions. To verify that Skyfall-GS learns a coherent, globally explorable 3D structure rather than overfitting to IDU-sampled viewpoints, we include free-flight trajectory renders along novel, smooth camera paths withheld during training. These renders confirm that our method produces a fully immersive and structurally consistent exploration experience, and enable direct perceptual comparison against baseline approaches and Google Earth Studio reference videos.
Appendix 0.B Implementation Details
0.B.1 Method Components
0.B.1.1 Codebase.
Our method extends the Mip-Splatting [111] codebase with custom modules for satellite imagery processing and our curriculum-based IDU refinement pipeline.
0.B.1.2 Pseudo camera depth supervision.
We sample cameras with varied azimuths and decreasing elevations, using random per-image embeddings. MoGe [90] provides scale-invariant depth estimation. We sample 24 views every 10 iterations, with look-at points $(x,y,z)$, where $x,y\sim\mathcal{N}(0,128)$ and $z=0$. Camera azimuths are uniformly sampled between 0 and $2\pi$, while elevation angles and radii linearly decrease from $80\degree$ to $45\degree$ and 300 to 250 units, respectively. Rendered RGB images ($I_{\text{RGB}}$) are $1024\times 1024$ pixels. We illustrate the 3DGS rendered RGB image $I_{\text{RGB}}$, scale-invariant depth $D_{\text{est}}$ estimated by MoGe [90] and depth from 3DGS $D_{\text{GS}}$ in Figure˜11.
@@ 0.B.1 Method Components
Figure 11: Pseudo-cam Depth Supervision. We use MoGe [90] to estimate the scale-invariant depth $D_{\text{est}}$ from the rendered RGB image $I_{\text{RGB}}$. The rightmost figures show the rasterized depth $D_{\text{GS}}$ from 3DGS.
0.B.1.3 3DGS reconstruction from satellite imagery.
@@
Our satellite-view optimization process runs for 30,000 iterations, with densification enabled between iterations 1,000 and 21,000. We modify several key parameters in the standard 3DGS implementation to address satellite imagery’s unique challenges. First, to prevent undesirable Gaussian elongation artifacts common with overhead views, we reduce the scaling learning rate from 0.005 to 0.001. Second, we implement pruning of Gaussians with maximum covariance exceeding 20 to eliminate floating artifacts. The loss function weights are set to $\lambda_{\text{D-SSIM}}=0.2$, $\lambda_{\text{op}}=10$, and $\lambda_{\text{depth}}=0.5$ for optimal reconstruction quality. For the GoogleEarth dataset, we disable the opacity regularization, since the dense training views are sufficient to eliminate floaters. For appearance modeling, we adopt the architecture from WildGaussians [35], implementing an appearance MLP with 2 hidden layers (128 neurons each) and ReLU activation functions. The per-image and per-Gaussian embedding dimensions are set to 32 and 24, respectively, with learning rates of 0.001, 0.005, and 0.0005 for per-image embeddings $e_{j}$, per-Gaussian embeddings $g_{i}$, and the appearance MLP $f$, respectively. The complete satellite-view training requires approximately 1 hour on a single NVIDIA RTX A6000 GPU.
0.B.1.4 FlowEdit-based refinement.
@@
We set FlowEdit noise parameters $n_{\text{min}}=4$ and $n_{\text{max}}=10$ to balance artifact removal with detail preservation. Our source prompt (“Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts.”) characterizes initial renders, while the target prompt (“Clear satellite image of an urban area with sharp buildings, smooth edges, natural lighting, and well-defined textures.”) guides refinement. These parameters were determined through experimentation, with lower noise values preserving more original structure but removing fewer artifacts, and higher values creating more significant changes but potentially altering underlying geometry. All other FlowEdit parameters use default values.
0.B.1.5 Curriculum-based refinement.
@@
Our IDU process comprises $N_{e}=5$ episodes of 10,000 iterations each, with densification through iteration 9,000. At the start of IDU, we randomly select and fix a single per-image appearance embedding $e_{j}$. Opacity regularization is disabled during IDU, as our curriculum naturally mitigates floating artifacts through multi-view consistency, enabling Gaussians to retain variable opacities beneficial for semi-transparent structures [33]. For DFC2019 [40] dataset, we utilize $N_{p}=9$ look-at points in a $3\times 3$ grid (512 units wide, centered at origin), with $N_{v}=6$ cameras per point and $N_{s}=2$ samples per view. Camera elevations decrease from $85^{\circ}$ to $45^{\circ}$ and radii from 300 to 250 units across episodes. For GoogleEarth [101] dataset, we utilize $N_{p}=16$ look-at point at origin, with $N_{v}=6$ cameras per point and $N_{s}=2$ samples per view. Camera elevations decrease from $85^{\circ}$ to $45^{\circ}$ and radius is fixed 600-unit across episodes. All training images are rendered at $2048\times 2048$ resolution. Our training strategy samples 75% from refined images and 25% from original satellite images, this sampling strategy makes sure that the final 3DGS scene faithfully follows the semantic and layout in the input satellite imagery. The complete synthesizing stage requires approximately 6 hours on a single NVIDIA RTX A6000 GPU.
0.B.2 Resource and Approximation Analysis
0.B.2.1 Rendering efficiency.
Our method achieves 11 FPS on the modest NVIDIA T4, significantly outperforming CityDreamer’s 0.18 FPS, which runs on the far more powerful NVIDIA A100 ($5\times$ the CUDA cores, $10\times$ the memory bandwidth). GaussianCity reaches comparable speeds (10.72 FPS) but requires the high-end A100. Furthermore, our fused representation enables real-time rendering at 40 FPS on consumer hardware (MacBook Air M2), demonstrating that our method enables high-quality 3D urban navigation without specialized computing resources.
0.B.2.2 Memory consumption.
We distinguish between peak memory and final memory. The peak memory usage reaches 46 GB during the synthesis stage, driven by the overhead of loading the diffusion model (FLUX.1) and temporary densification of Gaussians. However, the final training memory footprint is significantly lower (28.04 GB) as our method actively prunes redundant and low-opacity points. In terms of scene complexity, the refinement process densifies the scene by approximately 27%, increasing the Gaussian count from $\sim$1.65 million (reconstruction stage) to $\sim$2.1 million, specifically targeting the vertical facade geometry missing in the initial satellite reconstruction.
0.B.2.3 Validity of RPC to perspective approximation.
We adopt the methodology proposed in SatelliteSfM [115] to approximate the satellite linear pushbroom sensor as a perspective camera. This approximation relies on the “weak perspective” assumption, which holds valid when the satellite altitude ($Z$) is significantly larger than the depth variation within the scene ($\Delta Z$), i.e., $Z\gg\Delta Z$. Given that satellites orbit at distances of hundreds of kilometers while terrestrial depth variations are limited to a few hundred meters, the ratio $\Delta Z/Z$ remains negligible, allowing the geometry to converge to a perspective model. The approximation is achieved by generating a dense grid of 3D-2D correspondences using the rigorous RPC model and solving for a projection matrix $P$ via the Direct Linear Transformation (DLT) method, which is subsequently decomposed ($P=K[R|t]$) to recover camera parameters. Quantitative evaluations demonstrate that this process introduces negligible error: the average maximum forward projection error against the rigorous RPC model is only 0.126 pixels, and the difference in triangulated 3D points is typically less than 5 cm. Furthermore, this initialization allows Bundle Adjustment to achieve sub-pixel accuracy, with median reprojection errors recorded at 0.864 pixels, confirming the suitability of this approximation for high-fidelity 3D reconstruction.
Appendix 0.C Main Paper Experiments Detail & Results
0.C.1 Dataset Details
0.C.1.1 DFC2019 dataset.
@@ 0.C.1 Dataset Details
The number of training images and geographical coordinates for each AOI is provided in Table˜5. We also include four additional AOIs from Jacksonville to demonstrate our method’s robustness across varying scene characteristics. The number of training images and geographical coordinates for these additional AOIs is provided in Table˜6. These additional AOIs feature distinct characteristics: one contains a city hall building (JAX_164), another includes an American football stadium (JAX_175), while the remaining two exhibit other notable urban features (JAX_168 and JAX_264).
0.C.1.2 GoogleEarth dataset.
@@
The GoogleEarth dataset, introduced by CityDreamer [101], contains semantic maps, height fields and renders from Google Earth Studio [22] of New York City. This dataset is used to train the generative model in CityDreamer [101] and GaussianCity [103]. We pick four AOIs which contain diverse city elements, including complex architectures (004), squares (010), resident area (219) and riverside (336). However, original GES renders provided in GoogleEarth dataset are rendered from a lower elevation angle, which is not similar to satellite imagery. Therefore, for each AOI, we render 60 images from GES using an orbit trajectory with $80^{\circ}$ of elevation angle and 2219 meter of radius. These new renders serve as the input of our methods. The AOI ID, geographical coordinates, and the number of input images are detailed in Table˜7.
@@
Table 5: Number of training images and geographical coordinate per Area of Interest (AOI). These AOIs correspond to standard evaluation scenarios established by previous works, ensuring consistent and fair comparisons with existing baselines (e.g., Sat-NeRF [56]).
@@
Table 7: Number of training images and geographical coordinate per Area of Interest (AOI). We pick 4 AOIs from the GoogleEarth [101] dataset, ensuring fair comparisons with existing baselines (e.g., CityDreamer  [101] and GaussianCity [103])
@@
0.C.2 Evaluation Protocol
0.C.2.1 User study.
We asked participants three specific questions and instructed them to select one video that best addressed each question: